# Role of volatility mixing in wealth condensation transition

### 15 Apr 2026 v1

@@ II Model
where $c_{i}$ denotes the normalized wealth [1] of node $i$, $t$ is time, $W_{t,i}$ is a Wiener process of $i$ at $t$, $\beta_{i}=\beta_{\pm}=\beta\pm\Delta\beta$ is the volatility of node $i$, $J$ is the strength of interaction between connected nodes, $\mathbf{a}$ is the adjacency matrix for a given network with the element $a_{ij}$, $\langle{k}\rangle$ is the mean degree, and $N$ is the total number of nodes. Throughout this paper, Eq. (2) is interpreted in the Itô interpretation. In contrast to the homogeneous model, the volatility configuration $\bm{\beta}=(\beta_{1},\dots,\beta_{N})$ for a given network plays the role of a quenched disorder and also influences a wealth distribution.
@@
where $p$ denotes the overall probability of connection between nodes (link) in the SBM, and $q$ is the probability of connection between blocks (inter-block connection). For a given mean degree $\langle{k}\rangle$, the corresponding ER network has a connection probability $p_{0}=\langle{k}\rangle/(N-1)$, while the SBM has $p\approx 2p_{0}$ in the large limit of $N$, so the mean degree is independent of $q$. In particular, when $q=1/2$, the SBM is equivalent to an ER network in which every pair of nodes is connected with probability $p/2$. The assignment of $\beta_{+}$ and $\beta_{-}$ to the two blocks yields
@@ Appendix A Degree distributions for ER network and stochastic block model
In this appendix, we compare the network structures and degree distributions of the ER network and the stochastic block model (SBM). Let the total number of nodes be $N$ (assuming that it is even). In an ER network, the degree $k$ follows a binomial distribution,
@@
which reduces to $\mathbb{E}[k]\simeq Np/2$. in the large limit of $N$. Therefore, for fixed $k_{0}$ and $N\to\infty$, the following relationship is established.
@@
and the SBM reduces to two disconnected ER networks of size $N/2$, while preserving the same mean degree in the large limit of $N$. For $q=1/2$, the two binomial variables have the same success probability, and thus,
@@
For $0<q<1/2$, the SBM is not strictly equivalent to the ER network, since $k$ is the sum of two binomial variables with different probabilities. Nevertheless, its mean and variance converge to those of the corresponding ER network as $N\to\infty$. Therefore, for sufficiently large $N$, the degree distribution and the overall network structure of the SBM are nearly identical to those of an ER network with a mean degree $k_{0}$.
@@ C.1 Power-law test and size dependence
These observations suggest that in the small-$\Lambda$ regime, the truncated log-normal fit may reflect finite-size rounding of the tail sector rather than the asymptotic form itself. In particular, as $N$ increases, the tail extends to larger $c$ and the fitted log-normal tail becomes progressively flatter, while the effective exponent $\hat{\gamma}$ remains approximately stable. This trend is consistent with a crossover to the power-law scaling in the thermodynamic limit of $N\to\infty$. We therefore interpret $\hat{\gamma}$ as an effective exponent against scaling, while avoiding the claim that finite-$N$ samples are uniquely characterized by a pure power-law over the entire fitted range. Within this interpretation, the cases with $\hat{\gamma}\leq 2$ are consistent with condensation in the thermodynamic limit.

### 16 Apr 2026 v2

@@ II Model
where $c_{i}$ denotes the normalized wealth [1] of node $i$, $t$ is time, $W_{t,i}$ is a Wiener process of $i$ at $t$, $\beta_{i}=\beta_{\pm}=\beta\pm\Delta\beta$ is the volatility of node $i$, $J$ is the strength of interaction between connected nodes, $\mathbf{a}$ is the adjacency matrix for a given network with the element $a_{ij}$, $\langle{k}\rangle$ is the mean degree, and NN is the total number of nodes. Throughout this paper, Eq. (2) is interpreted in the Itô interpretation. In contrast to the homogeneous model, the volatility configuration $\bm{\beta}=(\beta_{1},\dots,\beta_{N})$ for a given network plays the role of a quenched disorder and also influences a wealth distribution.
@@
where $p$ denotes the overall probability of connection between nodes (link) in the SBM, and $q$ is the probability of connection between blocks (inter-block connection). For a given mean degree $\langle{k}\rangle$, the corresponding ER network has a connection probability $p_{0}=\langle{k}\rangle/(N-1)$, while the SBM has $p\approx 2p_{0}$ in the large limit of NN, so the mean degree is independent of $q$. In particular, when $q=1/2$, the SBM is equivalent to an ER network in which every pair of nodes is connected with probability $p/2$. The assignment of $\beta_{+}$ and $\beta_{-}$ to the two blocks yields
@@ Appendix A Degree distributions for ER network and stochastic block model
In this appendix, we compare the network structures and degree distributions of the ER network and the stochastic block model (SBM). Let the total number of nodes be NN (assuming that it is even). In an ER network, the degree $k$ follows a binomial distribution,
@@
which reduces to $\mathbb{E}[k]\simeq Np/2$. in the large limit of NN. Therefore, for fixed $k_{0}$ and $N\to\infty$, the following relationship is established.
@@
and the SBM reduces to two disconnected ER networks of size $N/2$, while preserving the same mean degree in the large limit of NN. For $q=1/2$, the two binomial variables have the same success probability, and thus,
@@
For $0<q<1/2$, the SBM is not strictly equivalent to the ER network, since $k$ is the sum of two binomial variables with different probabilities. Nevertheless, its mean and variance converge to those of the corresponding ER network as $N\to\infty$. Therefore, for sufficiently large NN, the degree distribution and the overall network structure of the SBM are nearly identical to those of an ER network with a mean degree $k_{0}$.
@@ C.1 Power-law test and size dependence
These observations suggest that in the small-$\Lambda$ regime, the truncated log-normal fit may reflect finite-size rounding of the tail sector rather than the asymptotic form itself. In particular, as NN increases, the tail extends to larger $c$ and the fitted log-normal tail becomes progressively flatter, while the effective exponent $\hat{\gamma}$ remains approximately stable. This trend is consistent with a crossover to the power-law scaling in the thermodynamic limit of $N\to\infty$. We therefore interpret $\hat{\gamma}$ as an effective exponent against scaling, while avoiding the claim that finite-NN samples are uniquely characterized by a pure power-law over the entire fitted range. Within this interpretation, the cases with $\hat{\gamma}\leq 2$ are consistent with condensation in the thermodynamic limit.